OpenStudy (anonymous):

FIND THE PRIME FACTORIZATION OF 4620. help please.

OpenStudy (kirbykirby):

Since your number ends in 0: it's easy to divide by 10: \(4620 \div 10=462\) and you know the prime factors of 10 are 5 and 2 (Keep 5 and 2 in mind) Now you have an even number, so divide by 2: \( 462 \div 2=231\) (keep 2 in mind) Now I notice that 231 is divisible by 3 since 2+3+1=6, a number divisible by 3 (this is one of the tricks to know which numbers are divisible by 3) \(231 \div 3 = 77\) (keep 3 in mind) Now hopefully you notice that you have double repeating digits, meaning this is a multiple of 11: \(77 \div 11 = 7\) (keep 11 and 7 in mind) So, your factors: \(2^2\times 5 \times 3 \times 7 \times 11\)

OpenStudy (anonymous):

2 * 2 * 3 * 5 * 7 * 11